OpenStudy (briensmarandache):

V=(1/3)pi*r^2*h, suppose r=h/2; which I don't understand why or how that is. And when I plug that in I get V=(1/3)pi*(h^3/4), but the book has (1/12)pi*h^3. I know its simple I'm just not seeing it.

OpenStudy (phi):

you got \[ V = \frac{1}{3} \pi \frac{h^3}{4} \] you can multiply fractions: top times top and bottom times bottom you get \[ V = \frac{1}{3} \pi \frac{h^3}{4} = \pi \frac{h^3}{12} \] which you can write a few different ways: \[ V = \frac{1}{12} \pi h^3 \] or \[ V = \frac{\pi h^3}{12} = \frac{\pi }{12}h^3\]

OpenStudy (phi):

suppose r=h/2 which is the same as make the diameter (2 * r) equal to h it means: pick a height for the cone. then make its diameter of the base the same length (which you can do) it looks roughly like this